OpenStudy (anonymous):

Find a quadratic function to model the values in the table. Predict the value of y for x = 6. x y 1 2 0 2 3 -10

OpenStudy (helder_edwin):

the general quadratic function is \[\large y=ax^2+bx+c \] replace the three points u have and u will get three linear equations. then solve the system

OpenStudy (anonymous):

I tried that but i'm still not getting it =/

OpenStudy (shamim):

2=a+b+c 2=c -10=9a+3b+c solve this 3 equation for a,b,c

OpenStudy (shamim):

tell me the values of a,b ,c

OpenStudy (whpalmer4):

Plugging in the values to get the equations: \[y = ax^2+bx+c\]\[\begin{array}{cc} x & y \\\hline\\ 0 & 2 \\ 1 & 2 \\ 3 & -10 \\ \end{array}\] \[2 = a (0)^2 + b(0) + c\]\[2 = c\] \[2 = a(1)^2 + b(1) + c\]\[2 = a+b+c\]\[2 = a + b + 2\]\[0 = a+b\] \[-10 = a(3)^2 + b(3) + c\]\[-10 = 9a + 3b + 2\]\[-12 = 9a+3b\] So you have a system of 3 equations in 3 unknowns, although we've already found one of them, giving us a system of 2 equations in 2 unknowns: \[~~~~~0 = ~~a+~~b\]\[-12 = 9a+3b\] The first equation implies that \(a = -b\), so we can substitute into the second equation; \[-12 = 9(-b) + 3b\]What do you get for \(b\) when you solve that equation? From there, find the value of \(a\) and write your full equation (remember that \(c = 2\)). Use it to find the value of \(y\) when \(x=6\).